# From Language to Action: A Review of Large Language Models as Autonomous Agents and Tool Users

### v1

@@ 10 Discussion
In memory mechanisms, the divide between localized and distributed memory is pronounced. Single-agent models optimize within-episode retention using episodic reprompting [58], compressed logs [116], or symbolic graphs [49], supporting isolated task continuity. Multi-agent memory systems, such as ProAgent [25], support collective planning through shared task memory and thread-level history, while MATRIX [98] serves as a multi-agent simulation environment for generating diverse interaction data, rather than managing shared histories or memory. Widely used strategies like context windows [68, 130] and dialogue history [97, 65] underpin both categories, while hybrid memory systems [70, 222] signal convergence toward unified memory architectures that bridge short- and long-term retention.
Table 10: Identified research gaps and potential research paths for advancing LLM-based agent systems.
Research Gaps Potential Research Paths
Lack of Verifiable Reasoning 1. Replace unstructured chain-of-thought with logically consistent and auditable reasoning frameworks.
2. Integrate symbolic AI approaches (e.g., BDI agents) for verifiable decision-making.
3. Ground reasoning in domain-specific processes for technical and scientific accuracy.
4. Incorporate fine-grained error detection and correction mechanisms during inference.
Limited Self-Improvement 1. Employ introspection-based reinforcement learning for continuous error identification and correction.
2. Use cooperative multi-agent evolution to share and refine learned strategies.
3. Maintain adaptive capabilities without requiring full-scale retraining.
4. Leverage reflective feedback loops to iteratively enhance performance.
Infrastructure Bottlenecks 1. Optimize inference latency through KV-caching and low-latency decoding pipelines.
2. Implement lightweight, modular architectures for resource-constrained settings.
3. Enable efficient on-device and streaming computation for real-time tasks.
Weak Multi-Agent Communication 1. Design adaptive dialogue protocols inspired by human social cognition.
2. Integrate contextual modeling for dynamic, multi-agent negotiation and coordination.
3. Improve real-time collaboration in competitive and cooperative environments.
Limited Context-Sensitive Collaboration 1. Develop proactive goal refinement methods for dynamic user needs.
2. Enhance co-creative interaction to anticipate and adapt to evolving task requirements.
3. Move beyond passive instruction-following toward proactive task engagement.
Insufficient Personalization 1. Build persistent, evolving user models based on long-term interactions.
2. Incorporate privacy-preserving memory architectures.
3. Adapt multi-modal interfaces dynamically to user preferences.
Vulnerability to Adversarial Triggers 1. Establish robust detection and prevention mechanisms against backdoor attacks triggered by contextual or environmental signals.
2. Implement adversarial robustness testing during training and deployment.
3. Continuously monitor and respond to unsafe emergent behaviors.
Inadequate Explainability 1. Embed explainability directly into the reasoning process.
2. Provide step-by-step justifications for decisions in human-readable form.
3. Tailor explanations to different levels of user expertise.
Incomplete Evaluation Frameworks 1. Establish standardized benchmarks for multi-agent reasoning in complex contexts.
2. Include negotiation-heavy, ambiguous, and competitive tasks in evaluation.
3. Measure adaptability, reliability, and fairness across diverse scenarios.
Neglected Tool Adaptation 1. Treat agent tools as learnable and optimizable parameters.
2. Automate prompt engineering for task-specific and tool-specific optimization.
3. Evaluate the effect of tool adaptation on performance across varied tasks.
@@
Prompting, fine-tuning, and memory augmentation. The transformation of general-purpose LLMs into autonomous agents is guided by three basic mechanisms: prompt engineering, fine-tuning and memory augmentation (RQ5). Prompt engineering is a widely adopted non-parametric technique that allows flexible, inference-time control over agent behavior. We observed that systems such as MechAgents [34] and AGENTVERSE [35] extensively employed structured prompts to simulate expert collaboration and role play. They even used complex human-like behaviors, such as trust modeling and cognitive biases. However, developers resort to fine-tuning in scenarios where prompt-based control proves insufficient, particularly in highly specialized domains. This parametric approach enables agents to internalize domain expertise and task-specific skills, as seen in ToolLLM for tool use specialization, RISE for self-correction abilities, and SOTOPIA-$\pi$ [97] for modeling social intelligence. Likewise, memory augmentation is an increasingly critical mechanism. Our review shows that external memory, via RAG, and internal episodic memory are essential to ground agents in real-world contexts and support long-term reasoning. For example, TRIAGEAGENT [38] utilizes medical knowledge retrieval to mitigate hallucination, while Reflexion [13] enables agents to adapt based on accumulated experience. Therefore, these mechanisms interpret modern LLM agents’ architectural and behavioral autonomy, highlighting that agentic behavior is largely scaffolded by design choices rather than innate model capabilities.
@@
Evaluation and benchmarks. Our analysis indicated that the evaluation of agents based on LLM (RQ6) requires a methodological change, from relying solely on static NLP benchmarks to employing dynamic process-oriented evaluations that better capture the multifaceted nature of agent behavior. We also found that a robust evaluation framework should incorporate final outputs and the full agentic process. For example, AGENTBOARD [90] advances the field by introducing "progress rate" metrics that reflect how agents navigate intermediate steps rather than just results by measuring incremental task completion. Similarly, RE-Bench [89] aligns agent performance with human expert benchmarks on complex research problems, reinforcing the importance of human-level baselines in agent evaluation.
@@
In addition to this, controlled simulations have been used to observe how agents respond to the presence of unreliable or adversarial collaborators [95], highlighting resilience as a critical component of agent intelligence. Domain-specific evaluations also emerged as essential for assessing real-world utility. For example, ChemCrow [56] enables agents to autonomously plan and execute chemical syntheses in scientific applications, providing concrete metrics grounded in scientific outputs. In the healthcare domain, tool-integrated agents have shown measurable improvements by significantly reducing clinical calculation errors [55].
@@ 11 Future directions
In this section, we outline the key areas where future work and critical research is needed. illustrates the key terms. Table 10 provides an organized summary of the existing research gaps and potential research pathways discussed in this section.
@@ 11.1 Towards verifiable reasoning and robust self-improvement
Although current agents based on LLM show impressive abilities, their reasoning processes frequently remain unclear and their capacity to learn from failure is limited. This lack of clarity and inflexibility prevents their use in high-stakes, safety-critical, or scientifically demanding areas. Therefore, a key future direction involves creating agents with clear, verifiable reasoning methods and strong, efficient self-improvement cycles.
@@
One major trend in improving reasoning is shifting from unstructured, free-form "chain-of-thought" to more organized and logically sound frameworks. Some methods ensure logical consistency from the start. For example, the Theorem of Thought (ToTh) models reasoning as a collaboration among agents that simulate abductive, deductive, and inductive inference. It organizes the output into a formal reasoning graph, which is then checked for coherence [49]. Other works ground reasoning in clear, domain-specific processes. MechAgents [34] utilizes a team of agents to tackle complex engineering challenges by writing, executing, and self-correcting physics-based simulation code. This verifies the reasoning of the agent against established scientific models. Similarly, combining LLMs with traditional symbolic AI approaches, such as Belief-Desire-Intention agents, offers a way to enable verifiable decision-making and improve safety [103]. An essential aspect of verifiability is the ability to spot and address flaws in the reasoning process itself. FG-PRM research [223] tackles this by creating a detailed classification of errors in mathematical reasoning and training a reward model to identify and penalize these specific errors in each step of reasoning, thus increasing the reliability of the final product.
Alongside verifiable reasoning is the goal of strong self-improvement, where agents learn from experience without needing constant, large-scale retraining. A key strategy is giving agents (such as Reflexion) the power to reflect on their performance and learn from their mistakes, using verbal reinforcement learning, creating a memory of these reflections that guides future attempts [13]. Some studies (e.g., RISE) also employ recursive introspection to fine-tune models through an ongoing process of self-correction, helping the agent improve its output over time [43].
@@
Multi-agent dynamics also serves as a robust method for promoting self-improvement. In the CORY framework, a model is split into pioneer and observer agents that evolve together through a cooperative reinforcement learning process. This improves resilience and prevents the policy failures often seen in traditional RL fine-tuning [42]. Other frameworks, such as COPPER [111], feature a special reflector agent trained to provide tailored constructive feedback to other agents, improving teamwork through a systematic self-improvement loop. A particularly innovative direction is to improve the agent by adjusting its tools rather than its core model weights. Some studies, such as AgentOptimizer [112], suggest treating agent tools as learnable parameters, allowing the agent system to adapt to tasks by developing better tools. Likewise, AVATAR [51] automates the optimization of prompts for tool use by employing a comparator module to reason over both positive and negative examples. These combined methods of structured reasoning, introspective learning, and adaptable capabilities are leading the way for the next generation of more reliable, transparent, and continuously improving LLM agents.
11.2 Toward scalable, adaptive, and collaborative LLM-based agent systems
@@
As LLM-based agents continue to evolve, a crucial direction for future research lies in enhancing their scalability, adaptability, and real-time operational capacity, especially for deployment in multimodal and domain-specific environments [75, 80, 88, 106]. Although previous developments have emphasized internal agent reasoning and learning mechanisms, future work should address infrastructure-level constraints, including memory usage, latency, and efficient inference through architectural techniques such as KV cache and optimized decoding pipelines [71, 88]. These improvements are critical for supporting agents in low-resource or time-sensitive scenarios, such as on-device processing or streaming interaction.
@@
Another key priority lies in enhancing multi-agent collaboration and the development of perceptually aware communication, which goes beyond individual agent reasoning. In frameworks such as AGENTVERSE [35], the ability to negotiate, coordinate, and respond to dynamic multiagent environments remains limited due to the absence of adaptive communication protocols and the lack of contextual modeling [89]. Future agent systems should adopt adaptive dialogue management strategies inspired by human communication and social cognition to enable agents to facilitate contextual awareness and intent inference among agents [39].
@@
Furthermore, as the range of agent applications expands, extending to gesture-based interfaces, EHR prediction, or collaborative games, continuous learning, coevolutionary training, and prompt refinement methods become essential [75]. These techniques allow agents to evolve and adapt alongside changing tasks, users, or environments. Simultaneously, development of strategic evaluation methods, including negotiation-heavy simulations like Lewis-style coordination games or Diplomacy, can facilitate deeper understanding of agent reasoning under complex, ambiguous, and competitive settings [89, 97].
@@
Likewise, ensuring robustness and interpretability remains a primary concern in the advancement of LLM-based agents. Beyond introspective reasoning and verification strategies, complementary efforts are required to build transparent evaluation pipelines that track agent behavior over time, particularly in unpredictability or emergent behavior settings. The acceleration of open-source advancements, led by models such as LLaMA-2-70B, provides an accessible, scalable, and reproducible foundation for these future explorations, which enable empowering researchers to collectively develop safe, effective, and socially aligned agentic systems [99, 45].
@@ 11.3 Deepening the human-agent symbiosis: personalization, proactivity, and trust
A critical trajectory for future research involves enhancing the autonomy, alignment, and practical deployability of LLM-based agents, particularly within complex, dynamic, open-ended, and high-stakes environments. With increased operational and decision-making freedom, these agents begin to display emergent behaviors: some beneficial, such as strategic cooperation; and others potentially harmful, such as shortcut-seeking actions that undermine long-term safety. Addressing this duality requires the building of sophisticated and robust alignment strategies that can mitigate adverse behaviors while preserving constructive autonomy that enables agents to operate efficiently and effectively in complex real-world scenarios.
Simultaneously, the strategic focus for LLM agents is increasingly shifting toward increasing human performance by creating collaborative, reliable, and individualized assistance rather than replacing human capabilities. Realizing this human-AI partnership necessitates calls for advances in intent recognition and co-creative interaction, where agents can move beyond passive instruction follow-up to engage in context-sensitive conversations that anticipate user needs and refine task goals interactively, an approach illustrated by studies like "Tell Me More" [66] and ReHAC [102].
Long-term personalization represents a critical research frontier in which future agents should build persistent and evolving user models through sustained interaction over time while navigating the technical and ethical implications of user data management around memory architecture and privacy, as demonstrated in efforts like PrivacyAsst [94]. Achieving such personalization at scale demands multi-modal capabilities, where agents can integrate and respond to inputs like text, speech, and gesture as well while aligning interface modalities to user-specific preferences.
Alongside these capabilities, trust forms the foundation for effective human-agent collaboration. This requires resilient safeguard frameworks to protect against vulnerabilities such as backdoor attacks triggered by environmental signals as identified by Yang et al. [93], and equally, to advance the development of transparent and explainable reasoning mechanisms that allow agents to articulate their decision-making processes in human-understandable terms, toward explainable-by-design systems such as BDI-integrated LLMs [103].
Real-world scalability is highly dependent on continuous model optimization, leveraging innovative techniques such as collaborative decoding, lightweight modular execution strategies, and efficient caching mechanisms. Collectively, these integrated research directions aim to establish the foundation for new generation intelligent agents that can go beyond technical proficiency to ethically aligned, emotionally intelligent, privacy-conscious, and deeply integrated into complex human workflows.
@@ 12 Conclusion
In this article, we present a comprehensive overview of LLM-based agents and the integration of tools within these systems. We examined how prompt engineering, fine-tuning, memory enhancement, and tool use contribute to the building of LLM agents by addressing seven focused research questions. We observed that single-agent systems prioritize autonomy and introspective decision making. However, multi-agent systems focus on coordination, role distribution, and collaborative planning. Crucially, this distinction becomes domain sensitive, with multiagent configurations demonstrating pronounced advantages in areas requiring social intelligence, cooperative problem solving, and high-stakes decision support, such as healthcare, scientific research, and complex engineering tasks. Integrating external tools, real-time data sources, and multimodal systems has become essential to enable LLM agents to perform tasks beyond the limitations of pre-trained models. In parallel, the evaluation of these agents is changing from static accuracy-based benchmarks to dynamic process-oriented methods that account for reasoning quality, adaptability, and task completion in real-world settings.
@@
Conflict of interests: On behalf of all authors, the corresponding author states that there is no conflict of interest. Funding: No external funding is available for this research.
@@
Author Contributions: Conceptualization Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Mohaimenul Azam Khan Raiaan; Methodology: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Resources and Literature Review: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Writing – Original Draft Preparation: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Md. Abdur Rahman, Mohaimenul Azam Khan Raiaan;
Validation: Sami Azam, Mohaimenul Azam Khan Raiaan, Md. Abdur Rahman;
Formal Analysis: Sadia Sultana Chowa, Sami Azam, Mohaimenul Azam Khan Raiaan, Md. Abdur Rahman;
Writing – Reviewing and Finalization: Mohaimenul Azam Khan Raiaan, Sami Azam;

### v2

@@ 10 Discussion
We observe a clear divide between localized and distributed memory. Single-agent models optimize within episode retention using episodic reprompting [56], compressed logs [112], or symbolic graphs [47], supporting isolated task continuity. Multi-agent models like ProAgent [25] provide shared task memory and thread-level history, while MATRIX [94] generates diverse interaction data as a multi-agent simulation environment, rather than managing shared memory. Context windows [65, 125] and dialogue history [93, 22] underpin both categories, while hybrid memory systems [67, 214] signal convergence toward unified architectures that bridge short/long-term retention.
@@
Prompting, fine-tuning, and memory augmentation. Our findings showed that agent autonomy is enabled by prompt engineering, fine-tuning, and memory augmentation (RQ5). Prompt engineering, a non-parametric technique, allows flexible, inference-time control over agent behavior, as seen in MechAgents [33] and AGENTVERSE [34], which used structured prompts to simulate complex interactions such as expert collaboration, role play, trust modeling, and cognitive biases. However, in special domains where prompting is insufficient, fine-tuning is employed as a parametric approach to internalize domain expertise and task-specific skills, as seen in ToolLLM [50] for tool use specialization, RISE for self-correction abilities, and SOTOPIA-$\pi$ [93] for modeling social intelligence. Likewise, memory augmentation through external RAG and internal episodic memory is essential to ground agents in real-world contexts and support long-term reasoning. For example, TRIAGEAGENT [37] utilizes medical knowledge retrieval to mitigate hallucination, while Reflexion [13] enables agents to adapt based on accumulated experience. Therefore, the architectural and behavioral autonomy of modern LLM agents is largely scaffolded by design choices rather than innate model capabilities.
@@
Evaluation and benchmarks. Our analysis indicated that evaluating LLM-based agents (RQ6) requires a methodological change, from static NLP benchmarks to dynamic process-oriented evaluations. We also found that a robust evaluation framework should incorporate final outputs and the full agentic process. For example, AGENTBOARD [23] introduces "progress rate" metrics that capture incremental task completion across intermediate steps. Similarly, RE-Bench [86] aligns agent performance with human expert benchmarks on complex research problems, reinforcing the importance of human-level baselines in agent evaluation.
@@
In addition to this, controlled simulations reveal handling unreliable or adversarial collaborators [91], highlighting resilience as a critical component of agent intelligence. Domain-specific evaluations also emerged as essential for assessing practicality. For example, ChemCrow [54] benchmarks autonomous chemical synthesis against metrics grounded in scientific outputs, while tool-integrated agents in healthcare significantly reduce clinical calculation errors [53].
@@ 11 Future directions
In this section, we outline the key areas where future work and critical research is needed (RQ7). Table S3, in the Supplementary Material, provides an organized summary of the existing research gaps and potential research pathways discussed in this section.
@@ 11.1 Towards verifiable reasoning and robust self-improvement
Although current agents based on LLMs show impressive abilities, their reasoning processes frequently remain unclear, and their capacity to learn from failure is limited. This lack of clarity and inflexibility limits their usage. Future work should develop agents with clear, verifiable reasoning methods and strong, efficient self-improvement cycles.
@@
One major trend in improving reasoning is shifting from unstructured, free-form "CoT" to more organized and logically sound frameworks. Some methods ensure logical consistency from the start. For example, the ToTh model’s reasoning is a multi-agent collaboration that simulates abductive, deductive, and inductive inference, assembling outputs into a formal reasoning graph, validating for coherence [47]. In future works, a similar approach could assist in detecting financial fraud or security breaches by producing a reasoning graph that will provide a transparent, stepwise logic trail that authorities can trace and validate. MechAgents [33] utilizes a team of agents for complex engineering challenges and self-correcting physics-based simulation code. This verifies the reasoning of the agent against established scientific models. Building on this strategy, future work may explore the potential drug-drug interactions by modeling their combined molecular effects. Combining LLMs with traditional symbolic AI approaches, such as Belief-Desire-Intention agents, enables verifiable decision-making and improves safety [99]. An essential aspect of verifiability is identifying and addressing flaws in the reasoning process. FG-PRM research [212] reliably tackles this by creating a classification of errors in mathematical reasoning and training a reward model to identify and penalize these errors.
@@
Alongside verifiable reasoning is the goal of strong self-improvement, where agents learn from experience without needing retraining. Through strategies like Reflexion, agents can reflect on their performance and learn from mistakes, using verbal RL, creating a memory of these reflections that guides future attempts [13]. Studies such as RISE employ recursive introspection to fine-tune models through an ongoing process of self-correction, to improve their output over time [29]. Subsequent work could apply these strategies to develop an agent for biomedical image segmentation that self-improves iteratively, minimizes reliance on continuous large-scale human annotation, and adapts to patient-specific image differences. Multi-agent dynamics also serves as a robust method for promoting self-improvement. In the CORY framework, a model is split into pioneer and observer agents that co-evolve through a cooperative RL process. This improves resilience and avoids policy failures common in traditional RL fine-tuning [41]. Other frameworks, such as COPPER [107], feature a special reflector agent that provides tailored, constructive feedback to other agents, improving teamwork through a systematic self-improvement loop. Some studies, such as AgentOptimizer [108], suggest treating agent tools as learnable parameters, enabling task adaptation through tool improvement. Likewise, AVATAR [49] automates the optimization of prompts for tool use by employing a comparator module to reason over both positive and negative examples. Integrating structured reasoning, introspective learning, and adaptive capabilities is driving LLM agents toward more reliability, transparency, and continual improvement.
11.2 Towards scalable, adaptive, and collaborative LLM-based agent systems
@@
Future research should enhance LLM agents’ scalability, adaptability, and real-time operational capacity, for deployment in multimodal and domain-specific environments [72, 77, 85, 102]. Beyond internal agent reasoning and learning mechanisms, future work should address infrastructure constraints, memory usage, and latency, while enabling efficient inference through architectural techniques such as KV caching and optimized decoders [68, 85]. These improvements are vital for low-resource or time-sensitive scenarios, such as on-device processing or streaming interaction.
@@
Another key priority lies in enhancing multi-agent collaboration and the development of perceptually aware communication. The absence of adaptive communication protocols and contextual modeling in frameworks such as AGENTVERSE [34] limits their ability in dynamic multi-agent environments [86]. Future agent systems should adopt human-inspired, socially grounded dialogue strategies to enhance collaboration capability, contextual awareness and intent inference among agents [38]. Such improvements will enable the deployment of lightweight agents on edge hardware for traffic systems, warehouses, and drones, facilitating inter-agent communication and collaboration.
@@
Furthermore, as the range of agent applications expands, extending to gesture-based interfaces, EHR prediction, collaborative games, continuous learning, coevolutionary training, and prompt refinement methods become essential [72]. Simultaneously, development of negotiation-heavy simulations like Lewis-style coordination games or Diplomacy can facilitate deeper understanding of agent reasoning under critical settings [86, 93].
@@
Likewise, ensuring robustness and interpretability remains a primary concern. Beyond introspective reasoning and verification, transparent evaluation pipelines are needed to track agent behavior over time, particularly under unpredictability and emergence. The acceleration of open-source advancements, led by models such as LLaMA-2-70B, provides an accessible, scalable, reproducible foundation, enabling researchers to collectively develop safe, effective, socially aligned agentic systems [95, 43].
@@ 11.3 Deepening the human-agent symbiosis: personalization, proactivity, and trust
Future research should enhance autonomy, alignment, and practical deployability of LLM-based agents, for complex, dynamic, open-ended, and high-stakes environments. Increased operational freedom produces emergent effects, from strategic cooperative gains to unsafe shortcut-seeking actions. Addressing this duality requires robust alignment strategies to mitigate adverse behaviors while preserving constructive autonomy to operate efficiently and effectively.
Simultaneously, the strategic focus is shifting toward human-centered augmentation, emphasizing collaboration, reliability, and personalization. Realizing this calls for advances in intent recognition and co-creative interaction, where agents can engage in context-sensitive conversations that anticipate user needs and refine task goals interactively, an approach illustrated by studies like "Tell Me More" [63] and ReHAC [98].
Long-term personalization represents a critical research frontier in which future agents should build persistent, evolving user models over time while addressing memory and privacy, as stated in studies like PrivacyAsst [90]. Achieving such personalization requires integrating multimodal capabilities while aligning interface modalities to user-specific preferences. Recently, personalized LLM-based agents have introduced complex ethical and societal risks, anthropomorphism, and over-trust. Their humanlike interactions create deception and encourage vulnerable populations, such as children or the elderly, to disclose sensitive information [215]. Since LLM-based agents lack moral agency, attributing accountability is conceptually flawed [216]. Beyond interface risks, agent architectures suffer from jailbreaking [217], and environmental signal-triggered backdoor attacks [89]. Jailbreaking refers to deliberate overriding of built-in safety measures of LLMs [218], and since agents need to handle multi-round dialogues and multiple sources of information, they become more prone to jailbreaking attacks. Effective human-agent collaboration requires both resilient safeguard frameworks against vulnerabilities and transparent, explainable reasoning [99].
Continuous model optimization across collaborative decoding, lightweight modular execution strategies, and efficient caching mechanisms drives real-world scalability and enables a new class of agents that go beyond technical proficiency to ethically aligned, emotionally aware, privacy-conscious, and deeply integrated in complex workflows.
@@ 12 Conclusion
In this article, we present a comprehensive overview of LLM-based agents and the integration of tools within these systems. We examined how prompt engineering, fine-tuning, memory enhancement, and tool use contribute to the building of LLM agents by addressing seven focused research questions. We observed that single-agent systems prioritize autonomy and introspective decision making. However, multi-agent systems focus on coordination, role distribution, and collaborative planning. Crucially, this distinction becomes domain sensitive, with multi-agent configurations demonstrating pronounced advantages in areas requiring social intelligence, cooperative problem solving, and high-stakes decision support, such as healthcare, scientific research, and complex engineering tasks. Integrating external tools, real-time data sources, and multimodal systems has become essential to enable LLM agents to perform tasks beyond the limitations of pre-trained models. In parallel, the evaluation of these agents is changing from static accuracy-based benchmarks to dynamic process-oriented methods that account for reasoning quality, adaptability, and task completion in real-world settings.
@@
Conflict of interests: On behalf of all authors, the corresponding author states that there is no conflict of interest.
Funding: No external funding is available for this research.
@@
Author Contributions: Conceptualization Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Mohaimenul Azam Khan Raiaan; Methodology: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Resources and Literature Review: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Writing – Original Draft Preparation: Sadia Sultana Chowa, Riasad Alvi, Subhey Sadi Rahman, Md Abdur Rahman, Mohaimenul Azam Khan Raiaan;
Validation: Sami Azam, Mohaimenul Azam Khan Raiaan, Md Abdur Rahman, Md Rafiqul Islam;
Formal Analysis: Sadia Sultana Chowa, Sami Azam, Mohaimenul Azam Khan Raiaan, Md Abdur Rahman;
Writing – Reviewing and Finalization: Mohaimenul Azam Khan Raiaan, Sami Azam, Md Rafiqul Islam, Mukhtar Hussain;